Page with summary of Runs taken

Summary of all runs taken

You can find
  • the production list here of all the taken runs.
  • the first list here of all the FIRST taken runs.

The status of data reprocessing is documented here

11 August Morning. Production Run

File size increased to 1 GB. Tigger config. 2,4,6,9 (prescaled factor 8)

Run Description date notes
/first_r/DATA/production_0205.lmd (Aladin current=400units=500A, Steerer=0.000V dipole=3.165V) Beam Rate 0.9 kHz 11 Aug 2011 10:05 Fixed Triva problem

file 205 closed for error in MIMOSA FIFO 6 missing.

/first_r/DATA/production_0206.lmd | (Aladin current=400units=500A, Steerer=0.000V dipole=3.165V) Beam Rate 0.8 kHz | 11 Aug 2011 1:35 | Fixed Triva problem |

File 206 open after restart MIMOSA inside the cave

10-11 August night/morning. Production Run. After Triva restart

Tigger config. 2,4,6,9 (prescaled factor 8) Appears the triva problem

Run Description date notes
/first_r/DATA/production_0191.lmd (Aladin current=400units=500A, Steerer=0.000V dipole=3.165V) Beam Rate 1.1 kHz 11 Ago 2011 1:35 Fixed Triva problem
/first_r/DATA/production_0192.lmd (Aladin current=400units=500A, Steerer=0.000V dipole=3.165V) Beam Rate 1.1 kHz 11 Ago 2011 1:35 Fixed Triva problem
/first_r/DATA/production_0193.lmd (Aladin current=400units=500A, Steerer=0.000V dipole=3.165V) Beam Rate 1.1 kHz 11 Ago 2011 1:35 Fixed Triva problem
/first_r/DATA/production_0194.lmd (Aladin current=400units=500A, Steerer=0.000V dipole=3.165V) Beam Rate 1.1 kHz 11 Ago 2011 1:35 Fixed Triva problem
/first_r/DATA/production_0195.lmd (Aladin current=400units=500A, Steerer=0.000V dipole=3.165V) Beam Rate 1.1 kHz 11 Ago 2011 1:35 Fixed Triva problem
/first_r/DATA/production_0196.lmd (Aladin current=400units=500A, Steerer=0.000V dipole=3.165V) Beam Rate 1.1 kHz 11 Ago 2011 1:35 Fixed Triva problem
/first_r/DATA/production_0197.lmd (Aladin current=400units=500A, Steerer=0.000V dipole=3.165V) Beam Rate 1.1 kHz 11 Ago 2011 1:35 Fixed Triva problem
/first_r/DATA/production_0198.lmd (Aladin current=400units=500A, Steerer=0.000V dipole=3.165V) Beam Rate 1.1 kHz 11 Ago 2011 1:35 Fixed Triva problem
/first_r/DATA/production_0199.lmd (Aladin current=400units=500A, Steerer=0.000V dipole=3.165V) Beam Rate 1.1 kHz 11 Ago 2011 1:35 Fixed Triva problem

Fixed a small problem with X server on DAQ computer.

Run Description date notes
/first_r/DATA/production_0200.lmd (Aladin current=400units=500A, Steerer=0.000V dipole=3.165V) Beam Rate 1.1 kHz 11 Ago 2011 1:35 Fixed Triva problem
/first_r/DATA/production_0201.lmd (Aladin current=400units=500A, Steerer=0.000V dipole=3.165V) Beam Rate 1.1 kHz 11 Ago 2011 1:35 Fixed Triva problem
/first_r/DATA/production_0202.lmd (Aladin current=400units=500A, Steerer=0.000V dipole=3.165V) Beam Rate 1.1 kHz 11 Ago 2011 1:35 Fixed Triva problem
/first_r/DATA/production_0203.lmd (Aladin current=400units=500A, Steerer=0.000V dipole=3.165V) Beam Rate 1.1 kHz 11 Ago 2011 1:35 Fixed Triva problem
/first_r/DATA/production_0204.lmd (Aladin current=400units=500A, Steerer=0.000V dipole=3.165V) Beam Rate 1.1 kHz 11 Ago 2011 1:35 Fixed Triva problem

10 August. First Production Run.

Tigger config. 2,4,6,9 (prescaled factor 3)

Run Description date notes
/first_r/DATA/production_0180.lmd (Aladin current=400units=500A, Steerer=0.000V dipole=3.165V) Beam Rate 1.1 kHz 10 Ago 2011 19:20 prod run : ok
/first_r/DATA/production_0181.lmd (Aladin current=400units=500A, Steerer=0.000V dipole=3.165V) Beam Rate 1.1 kHz 10 Ago 2011 19:27 prod run : ok
/first_r/DATA/production_0182.lmd (Aladin current=400units=500A, Steerer=0.000V dipole=3.165V) Beam Rate 1.1 kHz 10 Ago 2011 19:35 closed at 426MB MIMOSA FIFO 6 missing
/first_r/DATA/production_0183.lmd (Aladin current=400units=500A, Steerer=0.000V dipole=3.165V) Beam Rate 1.1 kHz 10 Ago 2011 19:35 no good
/first_r/DATA/production_0184.lmd (Aladin current=400units=500A, Steerer=0.000V dipole=3.165V) Beam Rate 1.1 kHz 10 Ago 2011 19:35 restarted MIMO seems ok

Fixed Mimo problem
Run Description date notes
/first_r/DATA/production_0185.lmd (Aladin current=400units=500A, Steerer=0.000V dipole=3.165V) Beam Rate 1.1 kHz 10 Ago 2011 19:35 OK
/first_r/DATA/production_0186.lmd (Aladin current=400units=500A, Steerer=0.000V dipole=3.165V) Beam Rate 1.1 kHz 10 Ago 2011 19:35 OK
/first_r/DATA/production_0187.lmd (Aladin current=400units=500A, Steerer=0.000V dipole=3.165V) Beam Rate 1.1 kHz 10 Ago 2011 19:35 OK

Appears the triva problem. Run 190 only with land cosmics trigger.
Run Description date notes
/first_r/DATA/production_0190.lmd (Aladin current=400units=500A, Steerer=0.000V dipole=3.165V) Beam Rate 1.1 kHz 11 Ago 2011 1:35 Triva problem

9-10 August Night. Alignment Campaign

Trigger: 2,4,6,9 After spending few hours on the high kentros rate problem with decided to set high threshiolds (120 mV) to keep the rate under control and the run stable. All the data are taken with kentros with high th. Beam monitor is with Th. at 17 mV.
  • After 25 min 1 error from MIMOSA shows up: *-R2F-37 :read_meb :Mimosa FIFO 6 missing file (137, 140)
  • We quote the current read on the monitor. Please remember that the current effective is the one you read multiplied by 1.25
  • Start Counter Channel 0 is saturating: needs to be re-equalized (lowering gain)!
  • 5:00 we realize that there is a problem with the tdct of slat ~48 in the TOF. The tdct is not working... We request an access to verify if the tof is alive or not: TOF is fine.... We need to understand what are the drawbacks of setting the beam position on a slat with tdct NOT working.

Batch processing of all runs has already been submitted. Output will show up in /SAT/s/s371/PRODUCTION_root/. You will find out 2 files:
  • _rec_vtx.root (containing ALL the decoded information)
  • _rec_histo_vtx.root (containing the Shifter histograms)
Please take a look at the data and provide feedback!

Run Description date
/first_r/DATA/align_0132.lmd , 133, 134, 135 & 136 Run for checking alignment. Started with: (Aladin current=386, Steerer V=0 dipole=3.165) Beam Rate 3.5 kHz. Kentros rate on spill <100 Hz. Rate of Mimosa errors: 30/120k evts . 10 Ago 2011 3:27

/first_r/DATA/align_0137.lmd , 138,139 Run for checking alignment. 'Base condition': (Aladin current=375, Steerer V=0 dipole=3.165) Beam Rate 3.5 kHz. Kentros rate on spill <100 Hz. Rate of Mimosa errors: 30/120k evts . 10 Ago 2011 4:20

/first_r/DATA/align_0140.lmd , 141, 142 Run for checking alignment. Conditions: (Aladin current=360, Steerer V=0 dipole=3.165) Beam Rate 3.5 kHz. Kentros rate on spill <100 Hz. Rate of Mimosa errors: 30/120k evts . 10 Ago 2011 4:20

/first_r/DATA/align_0143.lmd, 144 145 Run for checking alignment. Conditions: (Aladin current=400, Steerer V=0 dipole=3.165) Beam Rate 3.5 kHz. Kentros rate on spill <100 Hz. Rate of Mimosa errors: 30/120k evts . 10 Ago 2011 4:20

/first_r/DATA/align_0146.lmd, 147 148 Run for checking alignment. Conditions: (Aladin current=420, Steerer V=0 dipole=3.165) Beam Rate 3.5 kHz. Kentros rate on spill <100 Hz. Rate of Mimosa errors: 30/120k evts . 10 Ago 2011 4:20

/first_r/DATA/align_0149.lmd, 150, 151 Run for checking alignment. Conditions: (Aladin current=440, Steerer V=0 dipole=3.165) Beam Rate 3.5 kHz. Kentros rate on spill <100 Hz. Rate of Mimosa errors: 30/120k evts . 10 Ago 2011 4:20

From 0152 we use an aladin current of 400.

9 August Afternoon. Waiting for magnet to come back.

Run Description date
test_no_magnet_0127 Testing reconstruction with no magnet 8 Ago 2011 24:00
test_no_magnet_0128 Testing reconstruction with no magnet 8 Ago 2011 24:00
test_no_magnet_0129 Testing reconstruction with no magnet 8 Ago 2011 24:00
test_no_magnet_0130 Testing reconstruction with no magnet 8 Ago 2011 24:00

8-9 August Night. Kentros setup run

Run Description date
global_t_0081 run for Kentros pedestal and alignment Mimosa/BM 8 Ago 2011 24:00
global_t_0082 - global_0089 run for alignment Mimosa/BM (Aladin current=375, Steere V=0), Mimosa noisy 8 Ago 2011 24:00
global_0090 run for alignment Mimosa/BM (Aladin current=375, Steerer V=0) 8 Ago 2011 24:00
global_0091 run for alignment Mimosa/BM (Aladin current=375, Steerer V=1) 8 Ago 2011 24:00
global_0092 run for alignment Mimosa/BM (Aladin current=375, Steerer V=-1 dipole=3.137=default) 8 Ago 2011 24:00
global_0093 run for alignment Mimosa/BM (Aladin current=375, Steerer V=0 dipole=3.237) 8 Ago 2011 24:00
global_0094 run for alignment Mimosa/BM (Aladin current=375, Steerer V=0 dipole=3.037) 8 Ago 2011 24:00
global_0095 run for alignment Mimosa/BM (Aladin current=375, Steerer V=0 dipole=3.100) 8 Ago 2011 24:00
global_0101 run for alignment Mimosa/BM (Aladin current=375, Steerer V=0 dipole=3.100) 8 Ago 2011 24:00
global_0102 run for alignment Mimosa/BM (Aladin current=375, Steerer V=0 dipole=3.137=default) 8 Ago 2011 24:00
global_0103 run for alignment Mimosa/BM (Aladin current=375, Steerer V=0 dipole=3.12) 8 Ago 2011 24:00
global_0104 run for alignment Mimosa/BM (Aladin current=375, Steerer V=0 dipole=3.145) 8 Ago 2011 24:00
global_0105 run for alignment Mimosa/BM (Aladin current=375, Steerer V=0 dipole=3.155) 8 Ago 2011 24:00
global_0106 run for alignment Mimosa/BM (Aladin current=375, Steerer V=0 dipole=3.175) 8 Ago 2011 24:00
global_0107 run for alignment Mimosa/BM (Aladin current=375, Steerer V=0 dipole=3.165) 8 Ago 2011 24:00
global_0108 - global_t_0126 run for kentros (Aladin current=375, Steerer V=0 dipole=3.165) 8 Ago 2011 24:00

8 August Night. Veto Position Scan Campaign

Run Description date
global_t_0038 First Run taken to scan the Veto position. 430 A 7 Ago 2011 24:00
global_t_0039 Small Run, same conditions as before. 430 A 7 Ago 2011 24:00
global_t_0040 Scanning veto position. Aladin current to 450 A 7 Ago 2011 24:10
global_t_0041 Scanning veto position. Aladin current to 470 A. Should be completely off Veto. 7 Ago 2011 24:15
global_t_0042 Scanning veto position. Going other side. Aladin current to 410 A. 7 Ago 2011 24:15
global_t_0043 Scanning veto position. Going other side. Aladin current to 390 A. 7 Ago 2011 24:30
global_t_0044 Scanning veto position. Going other side. Aladin current to 400 A. 7 Ago 2011 24:35
global_t_0045 Scanning veto position. Going other side. Aladin current to 420 A. 7 Ago 2011 24:40
global_t_0046 Scanning veto position. Going other side. Aladin current to 380 A. 7 Ago 2011 24:46
global_t_0047 Scanning veto position. Going other side. Aladin current to 370 A. 7 Ago 2011 24:50
global_t_0048 Scanning veto position. Going other side. Aladin current to 360 A. 7 Ago 2011 24:50
global_t_0049 Scanning veto position. Going other side. Aladin current to 350 A. 7 Ago 2011 1:10
global_t_0050 Scanning veto position. Going other side. Aladin current to 340 A. 7 Ago 2011 1:18
global_t_0051 Seems that a good position is found. Aladin current to 375 A. 7 Ago 2011 1:34
global_t_0052 Steer Scan started. -1.2 V. Aladin @ 375A. 8 Ago 2011 2:02
global_t_0053 Steer Scan started. -1.3 V. Aladin @ 375A. 8 Ago 2011 2:07
global_t_0054 Steer Scan started. -1.4 V. Aladin @ 375A. 8 Ago 2011 2:15
global_t_0055 Steer Scan started. -1.5 V. Aladin @ 375A. 8 Ago 2011 2:22
global_t_0056 Steer Scan started. -1.35 V. Aladin @ 375A. 8 Ago 2011 2:30
global_t_0057 Steer Scan started. -1.7 V. Aladin @ 375A. 8 Ago 2011 2:41
global_t_0058 Steer Scan started. -0.5 V. Aladin @ 375A. 8 Ago 2011 2:47
global_t_0059 Steer Scan started. 0. V. Aladin @ 375A. 8 Ago 2011 2:57
global_t_0060 Steer Scan started. -2. V. Aladin @ 375A. 8 Ago 2011 3:02
global_t_0061 Steer Scan started. -4. V. Aladin @ 375A. 8 Ago 2011 3:10
global_t_0062 Steer Scan started. -10. V. Aladin @ 375A. 8 Ago 2011 3:20
global_t_0063 Steer Scan started. 4. V. Aladin @ 375A. 8 Ago 2011 3:27
global_t_0064 Steer Scan started. 2. V. Aladin @ 375A. 8 Ago 2011 3:27

8 August Night. Triggering on the VETO

global_t_0065 Steer Scan started. -1.1 V. Aladin @ 375A. 8 Ago 2011 3:27

-- AlessioSarti - 07 Aug 2011
Topic revision: r11 - 2011-08-12, AlessioSarti
 
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding GSI Wiki? Send feedback
Imprint (in German)
Privacy Policy (in German)